Consider two countries, A and B. Each country produces only two goods with 1,000 production units: Wine and Cheese. Country A can produce 400 bottles of wine or 200 pounds of cheese or any combination of two goods. At the same time, country B can produce 1200 bottles of wine or 300 pounds of cheese or any combination of two goods. Suppose that both countries maximize their utility when they consume wine and cheese in equal proportions.

1.

a) Absolute Advantage

b) Comparative Advantage

c) Trade-off of wine for cheese

2. Production and consumption situations in both countries before trade

3. Production and consumption situations in both countries after trade with the barter exchange rate of three bottles of wine for one pound of cheese.

All necessary elements for complete answer have been included. Please provide detailed answer in an organized manner.
Answer #1
Country Wine Cheese
Country A 400 :200
Country B 1200 300

a) Absolute advantage---Country B   has àbsolute advantage in producing both wine and cheese.

We can see the graph below-----

1200 WINE B (PPC) 400 A (PRC) 200 * 7 300 loo 150 200 CHEESE

As we observe from the graph that PPC curvr of country B is higher than that of country A in profuction of both wine and cheese., So, vountry B has àbsolute advantage.in both goods Production.

b) Comparative Advantage---

Country A has comparative advantage in producing cheese and country B has comparative advantage in producing Wine

Explanation-----

Comparative advantage in international trade is calculated on the basis of opportunity cost of producing one good as compared to another good.

The country with lower opportunity cost , will product that good in his country.

In the example--------

*

Country wine cheese Opportunity cost of producing cheese opportunity cost of producting wine Comparative Advantage
Country A 400 200 400/200=2 200/400=0•5 Wine ( lesser O.C)
Country B 1200 300 1200/300=4 300/1200=0•25 Cheese( lesser OC)

c) Trade off of wine for cheese-----

Country A should produce cheese and trade wine with country B and country B should produce wine and trade cheese with country A

2) Production and consumption situations in both countries without trade----

Country A was maximising its utility in producing 200 bottled of wine and 100 pounds of cheese

Country B was maximising its utility by producing 600 bottles of wine and 150 pounds of cheese

country A= 200W+100C

country B= 600W+ 150C

world output= 800W+ 250 C= 1050 ( total output)

C) Production and consumption situation after trade------+

* Country A= 0W+200C

country B= 1200w+0C

world trade= 1200 W+ 200 C= 1400( total output)

Barter exchange between both countries will be 6 bottles if wine for one pound of cheese( 1200/200=6)

Hope you understand.